# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v3

@@ 2 Related Work
Prior surveys on LLM-driven AMAS predominantly focus on technical dimensions such as system architectures, agent capabilities, and domain-specific applications of Agentic AI. However, there remains a gap in addressing governance and policy aspects. For example, reviews like [27] and [28] delve into simulated environments, inter-agent communication protocols, and performance benchmarks. These works mainly detail how agents coordinate or scale but fall short in examining adversarial threats, such as prompt injection attacks or agent collusion, lifecycle governance (e.g., continuous monitoring or regulatory compliance), or system-wide explainability, especially for emergent behaviors in complex systems.
More specialized surveys, such as [29], focus on natural language communication, while [30] details collaboration mechanisms, exploring interaction protocols and practical use cases in domains like question-answering systems, 5G networks, and Industry 5.0. Although [29] highlights security risks associated with communication channels, and [30] discusses deployment contexts, both remain narrowly centered on coordination strategies. There is a need for deeper analysis of robustness against adversarial manipulation (e.g., Byzantine failures), governance frameworks for multi-agent deployments, and explainability challenges at scale.
@@
There are some review articles that address TRiSM in AI, such as [32], tackle ethical and risk-related issues including alignment with human values, discuss fairness in decision-making [37], and defenses against privacy attacks (e.g., data leakage). However, these works typically target general ML systems and overlook the unique dynamics of LLM-based multi-agent systems. For instance, while they may advocate for privacy-preserving techniques like differential privacy, they do not address privacy risks in inter-agent data sharing or the distributed nature of LLM-driven systems. Moreover, critical TRiSM aspects [38] like governance (e.g., auditing autonomous agents) and explainability (e.g., tracing decisions across multiple agents) are often absent in recent discussions.
Current Study (Multi-Agent Systems and LLMs)
This review distinguishes itself by mapping TRiSM principles into the study of AMAS. Unlike prior studies (see Table 1), which offer limited insight into governance and policy, this work operationalizes TRiSM principles in practice. It provides actionable guidance for researchers, engineers, and policymakers, addressing critical gaps in both the technical AMAS literature and the broader responsible AI landscape.
3 Literature Review Methodology
To review the literature related to TRiSM as applied in Agentic AI systems, we adopted a structured methodology informed by best practices in systematic reviews [39, 40]. This section outlines our research objectives, data sources, inclusion criteria, and classification strategy.
@@ 3 Literature Review Methodology
(b) TRiSM pillars papers by topic.
@@
Figure 2: (a) Annual trend in Agentic AI research, showing a sharp rise post-2023. (b) Distribution of TRiSM-related publications across governance and safety topics.
Research Objectives The review was guided by the following four specific questions:
  • RQ1: What are the key trust, risk, and security challenges posed by LLM-enabled Agentic AI systems? How do current approaches establish and measure the trustworthiness of Agentic AI outputs and decisions?
  • RQ2: What technical and governance strategies have been proposed to address these challenges?
  • RQ3: How do existing approaches map onto the TRiSM pillars: explainability (including trustworthiness), model operations, security, and privacy?
  • RQ4: What gaps remain in current research, and what directions are promising for future work?
@@
Search Strategy To conduct a review on Agentic AI and TRiSM, we systematically searched major digital libraries, including IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar. The search encompassed publications from January 2022 to May 2025, with select seminal works prior to 2022 included to provide foundational context. We utilized Boolean search queries with an expanded set of keywords to capture the interdisciplinary intersection of LLM-based multi-agent systems and TRiSM considerations. The search terms were structured as follows:
@@
(“ Agentic AI” OR “multi-agent systems” OR “multi-agent LLMs” OR “AI agents” OR “autonomous agents” OR “intelligent agents” OR “collaborative AI” OR “distributed AI” OR “LLM-based agents” OR “agent coordination” OR “agent-based modeling” OR “swarm intelligence”)
@@
AND (“trust” OR “trustworthiness” OR “risk” OR “security” OR “safety” OR “governance” OR “oversight” OR “compliance” OR “explainability” OR “interpretability” OR “transparency” OR “privacy” OR “data protection” OR “robustness” OR “accountability” OR “ethical AI” OR “fairness” OR “adversarial attacks” OR “regulatory compliance” OR “bias mitigation” OR “system reliability”)
Inclusion and Exclusion Criteria We applied the following criteria during the screening process:
Inclusion Criteria:
  • Studies published between 2022 and 2025, to capture developments following the introduction of LLMs (mostly post-ChatGPT).
@@
  • Studies that explicitly discuss Agentic AI, Multi-Agents, or LLM-powered AI systems in the context of at least one TRiSM dimension: trust, risk, security, governance, explainability, or privacy.
@@
  • Peer-reviewed articles, preprints, or reports from credible sources (e.g., arXiv, NeurIPS, IEEE, ACM, Nature, or governmental standards bodies).
Exclusion Criteria:
@@
  • Papers focused solely on traditional, rule-based, or symbolic agents without any integration of LLMs or emergent coordination.
@@
  • Studies addressing only low-level ML components (e.g., training optimization, architecture design) without reference to agentic behavior or TRiSM concerns.
@@
  • Non-English papers or those lacking sufficient metadata (e.g., missing abstracts or publication details).
After screening titles and abstracts, we shortlisted 250 papers. A subsequent full-text review narrowed the selection to 180 primary studies that directly addressed one or more TRiSM pillars. In addition to peer-reviewed literature, we also incorporated relevant technical reports, blogs, and website links to contextualize and map the emerging landscape of Agentic AI within the TRiSM framework. An analysis on the literature is shown in Figure 2.
@@
Quality Assurance To ensure the reliability of included studies, we conducted a quality assessment adapted from established systematic review guidelines [41]. Each paper was evaluated against the following criteria: (1) Does the study clearly define its objectives in the context of agentic or LLM-based systems? (2) Are the methods, experiments, or architectural designs well-documented and reproducible? (3) Does the paper substantively engage with at least one of the trust, risk, security, explainability, or privacy aspects? (4) Does the study offer empirical, theoretical, or normative contributions relevant to Agentic AI governance? Each criterion was rated as low, medium, or high. Studies scoring “low” in more than two dimensions were excluded or flagged for contextual relevance only. Quality ratings were assigned independently by four reviewers, with discrepancies resolved through discussion.
@@ 4.1 Traditional AI Agents vs Agentic AI
Traditional agents operate through predefined rules [42], heuristic workflows [43, 44], or deterministic logic [45], and perform well in narrow, well-defined environments. Cognitive Agents [46], a subset of traditional agents, model human-like reasoning by integrating perception, memory, and decision-making, often using architectures like SOAR [47] or ACT-R [48]. However, they remain constrained by static knowledge bases and limited adaptability [46].
@@
This Agentic AI architecture supports longitudinal task execution and cross-domain generalization, overcoming the context fragmentation inherent in traditional approaches [52]. For example, systems like AutoGen [53] enable LLM-backed agents to decompose complex tasks, such as software development, through multi-step reasoning [54]. This capability of Agentic AI represents a significant shift from reactive to proactive problem-solving systems [49]. A comparison of traditional AI agents and Agentic AI systems is presented in Table 2.
4.2 LLM-Based Agentic AI System Architecture
Agentic Multi-Agent System (AMAS) represent an emerging paradigm in AI where multiple LLM-powered agents operate semi-autonomously, interact with external tools, and collaborate to achieve complex tasks. As illustrated in Figure 3, a typical AMAS architecture comprises several key components that together form a flexible yet highly dynamic ecosystem. At the core are multiple LLM-Based Agents, each capable of reasoning, planning, and tool invocation [55]. These agents access a Shared Toolchain Interface [56] to execute code, perform searches, or interact with domain-specific APIs.
@@ 4.2 LLM-Based Agentic AI System Architecture
Communication and coordination are facilitated through a Communication Middleware [57] that allows agents to share goals, observations, or intermediate results. A Task Manager or Orchestrator [58] governs high-level planning and delegates subtasks to agents based on their roles or specializations. Agents can read from and write to a World Model or Shared Memory [27], which stores contextual knowledge, system state, or evolving task data. Human oversight is supported through a Human-in-the-Loop Interface [44], enabling users to prompt, correct, or halt agent behavior.
@@
Table 3: Comparison of representative LLM-based Agentic AI frameworks across key design axes.
Framework (Year) Core LLM Planning Memory Tool Use Notable Features
@@
Table  3 presents representative systems and shows how each one maps these five design axes (LLM core, planning/ reasoning, memory, tool-use, environment interface) into practice. For example, AutoGPT pairs a GPT-4 core with a self-looped chain-of-thought planner and a vector-database memory to serve as both tool layer and environment interface.
5 Threats and Risks in LLM-based Agentic Multi-Agent Systems (AMAS)
@@ 5.1 Threats in Agentic AI
Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures [87]. Below, we discuss some of these major threats.
@@
  • Agent orchestration. This risk involves central or distributed control mechanisms for role assignment and workflow mediation. A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT and AutoGen . These orchestration vulnerabilities differentiate Agentic AI systems from conventional stateless, single-threaded LLM deployments.
@@ 5.2 Taxonomy of Risks
5.3 Real-World Examples
@@
Next, we discuss the AI TRiSM prism framework and map its core principles to a AMAS.
6 The TRiSM Framework for Agentic AI
@@ 6.1 Overview of AI TRiSM and Its Core Components
Table 4: TRiSM pillars in Agentic AI systems with aligned governance references.
Pillar Primary Objective Key Practices & Agentic-LLM Considerations Exemplar Standards / Metrics
Explainability Transparent multi-agent decisions SHAP, LIME, counterfactuals; reasoning-trace logs; cross-agent dependency mapping NIST AI RMF “Explain” [99]
ISO/IEC 24029-1 [100] , ISO/IEC 42005:2025 [101]
Metrics: fidelity, sufficiency, human-trust score
ModelOps Lifecycle governance of agents, models, orchestration Version-controlled prompts; CI/CD w/ multi-agent sims; drift monitoring; rollback policies ISO/IEC 24029-1 [100] , ISO/IEC 42005:2025 [101]
SRE MTTR/MTTD; $\Delta$-accuracy (model drift); pipeline pass-rate
Application Security Prevent prompt injection, spoofing, lateral exploits Prompt sanitation; least-privilege scopes; encrypted comms; sandboxed tool calls; anomaly detectors OWASP Top-10 for LLM Apps [102]
Jailbreak-success%; mean-time-to-detect; exploit CVSS
Model Privacy Safeguard sensitive data and shared memory Differential-privacy training; data minimization; homomorphic encryption; secure enclaves; audit logs GDPR Art. 25 [103]
HIPAA §164 [104]
Metrics: $\epsilon$-DP budget, leakage rate, access-audit pass-rate
@@
Originally highlighted in industry guidelines for AI governance [105], TRiSM provides a structured approach to managing the unique challenges of LLM-based “Agentic AI” systems. The TRiSM framework addresses trust concerns in Agentic AI by focusing on four key pillars: Explainability, Model Operations (ModelOps), Application Security, and Model Privacy [59, 23]. Figure 5 illustrates how these four TRiSM pillars map onto the core architectural layers of an LLM-based Agentic AI system.
@@
Figure 5: TRiSM-Aligned Architecture for LLM-Driven Agentic AI Systems. The system is organised into four vertical layers: Trust & Explainability, ModelOps, Planning + Reasoning, and Model Privacy, each supported by a cross-cutting Monitoring & Governance Plane that enforces policies, logging, and compliance triggers. External tools are accessed through a single Secure API gateway, while the user interface feeds back into both the Trust layer and the governance plane to enable end-to-end oversight.

### v4

@@ 2.3 Survey Landscape
Prior surveys on LLM-driven AMAS predominantly address system architectures, agent capabilities, and domain-specific applications of agentic systems. For example, [14, 15] examine simulated environments, inter-agent communication protocols, and performance benchmarks. These works clarify how agents coordinate and scale but give limited treatment to adversarial threats (e.g., prompt injection, collusion), lifecycle governance (e.g., continuous monitoring, audit, regulatory compliance), and system-wide explainability, especially for emergent behaviors in complex settings.
More specialized surveys, such as [16], focus on natural-language communication, while [17] details collaboration mechanisms and practical use cases in domains like question answering, 5G networks, and Industry 5.0. Although [16] highlights security risks associated with communication channels and [17] discusses deployment contexts, both remain narrowly centered on coordination strategies. Deeper analysis is needed on robustness to adversarial manipulation (e.g., Byzantine failures), governance frameworks for multi-agent deployments, and explainability at scale.
@@
Review articles addressing TRiSM-related themes (e.g., [11]; see also fairness-focused work [41]) discuss alignment with human values, fairness in decision-making, and defenses against privacy attacks (e.g., data leakage). However, these surveys typically target general ML systems and overlook dynamics unique to LLM-based multi-agent settings. For instance, while they advocate privacy-preserving techniques such as differential privacy, they rarely address privacy risks stemming from inter-agent data sharing or the distributed nature of LLM-driven workflows. Moreover, critical TRiSM aspects—such as governance (e.g., auditing autonomous agents) and explainability (e.g., tracing decisions across multiple agents), are often under-specified [42].
2.3.1 Positioning of this study
This review maps TRiSM principles to LLM-based AMAS, operationalizing governance, explainability, security, privacy, and lifecycle controls for multi-agent settings. Relative to prior surveys (Table 1), we foreground system-level risks (e.g., prompt infection, impersonation, memory poisoning), propose evaluation dimensions and composite measures for coordination and tool use, and offer actionable guidance for researchers, engineers, and policymakers.
2.4 Literature Review Methodology
We conducted a structured literature review to synthesize AI TRiSM principles for AMAS. Guided by established systematic-review practices [43, 44], this section states our research objectives and describes the data sources and search strategy, the inclusion/exclusion criteria and screening procedure, and the classification schema used for analysis.
@@ 2.4 Literature Review Methodology
(b) Papers by TRiSM pillar.
(c) Topic trends in Agentic AI (agents, security, explainability).
(d) Bibliography entries per year (1987–2025).
@@
Figure 2: Bibliographic analysis of Agentic AI research. (a) Annual trend in Agentic AI publications, showing a sharp rise post-2023. (b) Distribution of TRiSM-related publications across governance and safety topics. (c) Topic-specific trends in agents, security, and explainability. (d) Overall growth in bibliographic entries, highlighting acceleration since 2020.
Research questions (RQs)
The review was guided by the following three questions:
  • RQ1: What trust, risk, and security challenges are specific to LLM-based AMAS?
  • RQ2: Which governance, explainability, security, privacy, and lifecycle (TRiSM) controls are proposed or adapted for AMAS?
  • RQ3: How are AMAS evaluated (datasets, tasks, metrics), and where are the gaps?
Search strategy
@@
We systematically searched major digital libraries: IEEE Xplore, ACM Digital Library, SpringerLink, arXiv, ScienceDirect, and Google Scholar. The search covered publications from January 2022 to May 2025, with select seminal works prior to 2022 included for foundational context. We used Boolean queries and expanded keywords to capture the intersection of LLM-based multi-agent systems and TRiSM considerations. Example query template:
@@
("Agentic AI" OR "multi-agent systems" OR "multi-agent LLMs" OR "AI agents" OR "autonomous agents" OR "intelligent agents" OR "collaborative AI" OR "distributed AI" OR "LLM-based agents" OR "agent coordination" OR "agent-based modeling" OR "swarm intelligence")
AND
@@
("trust" OR "trustworthiness" OR "risk" OR "security" OR "safety" OR "governance" OR "oversight" OR "compliance" OR "explainability" OR "interpretability" OR "transparency" OR "privacy" OR "data protection" OR "robustness" OR "accountability" OR "ethical AI" OR "fairness" OR "adversarial attacks" OR "regulatory compliance" OR "bias mitigation" OR "system reliability")
Screening and eligibility
Two-stage screening (title/abstract, then full-text) used these criteria. Inclusion:
  • Studies published 2022–2025 (capturing the post-2022 LLM wave), plus select pre-2022 foundational work.
@@
  • Studies explicitly discussing Agentic AI, multi-agent, or LLM-powered systems in the context of at least one TRiSM dimension: trust, risk, security, governance, explainability, or privacy.
@@
  • Peer-reviewed articles and widely cited preprints or reports from credible sources (e.g., NeurIPS, IEEE, ACM, Nature, governmental/standards bodies, arXiv).
Exclusion:
@@
  • Papers focused solely on traditional, rule-based, or symbolic agents without LLM integration or emergent coordination.
@@
  • Studies addressing only low-level ML components (e.g., training optimization, architecture design) with no agentic behavior or TRiSM concerns.
@@
  • Non-English papers or those lacking sufficient metadata (e.g., missing abstract or publication details).
For empirical papers, we assessed clarity of threat models, reproducibility (code/data), evaluation rigor (baselines, ablations), and external validity. For surveys, we assessed coverage breadth, structure, and treatment of risks/governance. After screening titles and abstracts, we shortlisted 250 papers; full-text review yielded 180 primary studies addressing one or more TRiSM pillars. We also incorporated relevant technical reports and authoritative web resources to contextualize the rapidly evolving Agentic-AI/TRiSM landscape. A bibliometric summary appears in Figure 2.
Quality assurance
@@
To ensure reliability, we adapted quality criteria from established guidelines [45]. Each paper was evaluated on: (1) clear objectives for agentic or LLM-based systems; (2) well-documented methods/experiments or architecture (reproducibility where applicable); (3) substantive engagement with at least one of trust, risk, security, explainability, or privacy; (4) empirical, theoretical, or normative contributions relevant to Agentic-AI governance. Each criterion was rated low, medium, or high; studies rated low on more than two criteria were excluded or flagged as contextual only. Quality ratings were assigned independently by four reviewers; discrepancies were resolved via discussion (with inter-rater agreement checked on a subset).
@@ 3.1 Traditional AI Agents vs Agentic AI
Traditional agents operate through predefined rules [46], heuristic workflows [47], or deterministic logic [48], and perform well in narrow, well-defined environments. Cognitive Agents [49], model human-like cognitive functions, such as integrating perception, memory, and decision-making, often using cognitive systems architectures like [50] or ACT-R [51]. According to a recent work [52], authors distinguish between 4 types of knowledge: Concepts, Skills, Processes, Motives for cognitive agents. A key difference between Agentic AI and such cognitive architecture approaches is that the latter were mostly applied to robot planning and control (so applies to ‘physical operational environments’) but Agentic AI applies more broadly, also to non-physical operational environments (which could be made to operate entities in physical also environments).
@@
This Agentic AI architecture supports longitudinal task execution and cross-domain generalization, overcoming the context fragmentation inherent in traditional approaches [56]. For example, systems like AutoGen [57] enable LLM-backed agents to decompose complex tasks, such as software development, through multi-step reasoning [58]. This capability of Agentic AI represents a significant shift from reactive to proactive problem-solving systems [53]. A comparison of traditional AI agents and Agentic AI systems is presented in Table 2. For the scope of this work, we rely on LLM-based Agentic AI systems.
3.2 Multi-Agentic AI System (AMAS) Architecture
AMAS represents an emerging paradigm in AI where multiple LLM-powered agents operate semi-autonomously, interact with external tools, and collaborate to achieve complex tasks. As illustrated in Figure 3, a typical AMAS architecture comprises several key components that together form a flexible yet highly dynamic ecosystem. At the core are multiple LLM-Based Agents, each capable of reasoning, planning, and tool invocation [59]. These agents access a Shared Toolchain Interface [60] to execute code, perform searches, or interact with domain-specific APIs.
@@ 3.2 Multi-Agentic AI System (AMAS) Architecture
As shown in Figure 3), communication and coordination are facilitated through a Communication Middleware [61] that allows agents to share goals, observations, or intermediate results. A Task Manager or Orchestrator [62] governs high-level planning and delegates subtasks to agents based on their roles or specializations. Agents can read from and write to a World Model or Shared Memory [14], which stores contextual knowledge, system state, or evolving task data. Human oversight is supported through a Human-in-the-Loop Interface [63], enabling users to prompt, correct, or halt agent behavior.
@@
Table 3: Comparison of representative LLM-based Agentic AI frameworks across key design axes. Acronyms: CoT = Chain-of-Thought reasoning, FSM = Finite State Machine, PDDL = Planning Domain Definition Language, DAG = Directed Acyclic Graph, DB = Database, REPL = Read–Eval–Print Loop, API = Application Programming Interface, KV = Key–Value store, CLI = Command Line Interface, SDK = Software Development Kit, MCP = Model Context Protocol.
Framework Core LLM Planning Memory Tool Use Notable Features
@@
Semantic Kernel [91] Model-agnostic Agent planning framework Memory & context mgmt Skill-based plugins, OpenAPI Enterprise focus, .NET integration, security, vector-DB support
OpenAI Swarm [92] GPT-4o Agents + handoffs loop Stateless (no LT memory) Python functions, agent handoffs Minimalist, controllable, educational multi-agent orchestration
OpenAI Agents SDK [93] Provider-agnostic Agent loops, handoffs, deterministic flows Sessions manage conversation history Tools + guardrails, tracing Multi-agent workflows, safety guardrails, extensibility
Strands Agents [94] Model-agnostic Agent loop with multi-agent & streaming N/A Python tools, hot-reloading, MCP tools Lightweight SDK, scalable autonomous workflows
LlamaIndex Agents [95] Model-agnostic Task breakdown + planning Task memory module External tools & parameters Pre-built agent/tool architectures, custom workflows
@@
Table  LABEL:tab:agentic_architectures presents representative systems and shows how each one maps these five design axes (LLM core, planning/ reasoning, memory, tool-use, environment interface) into practice.
4 Threats and Risks in LLM-based Agentic Multi-Agent Systems
@@ 4.1 Threats in Agentic AI
Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures. Below, we discuss some of these major threats.
@@
  • Agent orchestration. This risk involves central or distributed control mechanisms for role assignment and workflow mediation. A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT [53] and AutoGen [98] . These orchestration vulnerabilities differentiate Agentic AI systems from conventional stateless, single-threaded LLM deployments.
@@ 4.2 Taxonomy of Risks
Real-World Examples
@@
Next, we discuss the AI TRiSM framework and map its core principles to a AMAS and how it can mitigate these aforementioned risks.
5 The TRiSM Framework for Agentic Multi-Agent Systems
@@ 5.1 Overview of AI TRiSM and Its Core Components
Originally highlighted in industry guidelines for AI governance [110], TRiSM provides a structured approach to managing the unique challenges of LLM-based “Agentic AI” systems. The TRiSM framework addresses trust concerns in Agentic AI by focusing on key pillars, which are: Explainability, ModelOps, Application Security, Model Privacy and Goverance [64, 13]. Figure 5 illustrates how these four TRiSM pillars map onto the core architectural layers of an LLM-based Agentic AI system.
@@
Figure 5: TRiSM-aligned architecture for Agentic AI systems. The system is organized into four layers: Trust and Explainability (e.g., reasoning traces, user feedback), ModelOps (prompt versioning, CI/CD, orchestration), Planning and Reasoning (multi-agent CoT), and Model Privacy (encrypted memory, differential privacy, audit logs). A cross-cutting Governance Plane enforces oversight via policies and logging. External tools are accessed through a secure API gateway. The user interface connects back to the trust and governance layers for human-in-the-loop (HITL) oversight. Abbreviations used: TRiSM = Trust, Risk, and Security Management; CoT = Chain-of-Thought; CI/CD = Continuous Integration and Deployment; DP = Differential Privacy.
Table 4: TRiSM pillars for Agentic (LLM-based) multi-agent systems: core controls, techniques, risks, evaluation facets, example systems, and governing frameworks.
Pillar Core controls (keywords) Techniques / patterns / artifacts Key risks mitigated Evaluation facets & example metrics Example systems / patterns Primary standards / frameworks
Explainability / Trust CoT logging; layered-CoT; inter-agent traceability; explainer agent; role-based interpretability; prompt audit trails; attention viz; decision provenance Layered-CoT [111]; LIME/SHAP [112, 113]; decision-provenance graphs [114, 115]; multi-agent SHAP [116]; attention maps [117, 118]; role-based interpretability [119]; RAG-linked justifications [120]; prompt/action logs [121, 122] Opaque chains; unverifiable reasoning; hidden inter-agent effects; spurious correlations; user miscalibration Trustworthiness: calibration/consistency; Explainability: fidelity, stability; User-centric: task utility, cognitive load; Coordination: cross-agent attribution; Composite: multi-facet scorecards [123] SciAgent [120]; MetaGPT [124]; ChatDev [30]; AutoGen [57] EU AI Act Arts 13–14 [125]; NIST AI RMF / GAI profile [3, 126]; OECD AI [127]; ISO/IEC 24029-1 [32]
ModelOps (Lifecycle) Versioning; lineage; CI/CD safety gates; rollout/rollback; multi-agent simulation; hierarchical monitoring; usage/cost SLOs; ITSM/IAM integration Prompt & agent-config versioning; pre-deploy safety tests; regression/drift checks; sandbox simulation; hierarchical monitors (agent/group/system) [128] Silent regressions; drift; bias re-introduction; unsafe changes; cost blowouts; observability gaps Trustworthiness: drift/robustness; User-centric: task success, latency; Composite: risk & cost scorecards; incident MTTR/MTTD MegaAgent (hierarchical monitoring) [128] ISO/IEC 42001 (AIMS) [ISO_IEC_42001_2023]; ISO/IEC 42005 [129]; ISO/IEC 23894 [iso23894 or [130]]; NIST AI RMF / GAI profile [3, 126]; ModelOps [131, 132]
Application Security Prompt hygiene; secure prefixes; sandboxed tools; allowlisted actions; plan-then-execute; least privilege (RBAC); authN/authZ; anomaly detection; red-teaming Prompt-injection patterns [133, 58]; plan-then-execute; action selector; tool isolation; output filters; cross-agent cross-checks; adversarial training [134]; HITL holds Prompt injection; tool abuse; data exfiltration; cross-modal injection; lateral movement; jailbreaks Trustworthiness: robustness/adversarial resilience; Coordination: agent-of-agent validation; Composite: residual risk index; incident rates HITL review gates [135]; enterprise dashboards (override/stop) [135]; ChemCrow safety pauses [136] OWASP LLM Top-10 [137, 138]; OECD Robustness [127]
Model Privacy & Data Protection DP; anonymization/pseudonymization; minimization; MPC; HE/FHE; TEEs (SGX/SEV/TrustZone); encryption; access logs; retention/consent DP budgets [139]; k-anonymity [140]; MPC/PUMA [puma2024 or [141]]; HE/FHE pipelines [142]; TEEs [143]; memory scoping; PII detectors Memorization; leakage via shared memory/messages; unauthorized access; weak consent/retention Trustworthiness: privacy loss; User-centric: data-minimization utility; Composite: privacy-risk score; audit pass rate Encrypted memories; scoped retrieval; consented logs GDPR Art 25/5/35 [34]; CCPA/CPRA [144]; HIPAA §164 [35]; ISO/IEC 27001/27701 [iso27001, iso27701]
Governance Human oversight; accountability; auditability; DPIA; policy/change control; role clarity; transparency notices; logging/traceability Oversight playbooks; two-person verification for high-risk; role-oriented modularity; audit trails; DPIA templates; policy gates Unclear responsibility; weak oversight; non-compliance; opaque decisions User-centric: oversight efficacy; Coordination: traceability across agents; Composite: compliance scorecards HITL confirm/override/stop; role-based review boards EU AI Act Art 14 [125]; NIST AI RMF [3]; OECD AI [127]; ISO/IEC 42001/42005 [align keys]; ISO/IEC 23894 [130]